# PoU: Proof-of-Use to Counter Tool-Call Hacking in DeepResearch Agents

### v1

@@ Abstract.
Retrieval-augmented generation (RAG) agents—such as the recent DeepResearch-style systems—extend large language models (LLMs) with autonomous information-seeking capabilities via external tools. While reinforcement learning (RL) has enabled impressive multi-step reasoning performance, we reveal a previously overlooked failure mode, termed Tool-Call Hacking: agents inflate reward signals by issuing superficially correct tool calls without genuinely using the retrieved evidence. This leads to (i) mode collapse into repetitive reliance on a single source and (ii) spurious grounding, where answers are weakly supported by cited content. To address this, we propose Proof-of-Use (PoU)—an evidence-grounded RL framework that enforces verifiable causal links between retrieved evidence, reasoning traces, and final answers. PoU operationalizes this through a unified step-wise contract combining syntactic citation validation, perturbation-based sensitivity rewards, and answer–evidence alignment objectives, ensuring that tool usage remains both interpretable and functionally grounded. Across seven QA benchmarks covering in-domain, out-of-domain, and two out-of-tool-distribution tasks, PoU consistently outperforms strong DeepResearch baselines in factual accuracy, evidence faithfulness, and tool-routing balance. These findings highlight the necessity of grounding RL-trained agents not merely in task outcomes, but in the causal use of retrieved information—offering a principled path toward trustworthy retrieval-augmented reasoning.
@@ 1. Introduction
Retrieval-augmented generation (RAG) agents represent a recent line of work that integrates external tools (e.g., web search)(Jin et al., 2025; Song et al., 2025; Zheng et al., 2025) into the reasoning process. By combining parametric knowledge with retrieved evidence, these agents have substantially advanced complex question answering and research assistance. However, in many knowledge-intensive settings, a single external knowledge source is often insufficient (Ma et al., 2025). This motivates scaling RAG to multi-source environments. In such settings, diverse knowledge repositories are encapsulated as callable interfaces—commonly referred to as “tools” in agent-based frameworks. Here, a “tool” denotes a retrieval interface—e.g., web search, knowledge graph, domain databases or private corpora—each with distinct coverage and reliability. The multi-source environment presents a compounded challenge: not only must agents maintain calibrated intermediate beliefs and ground their conclusions in retrieved evidence, but they must also plan strategic tool routing—selecting which tools to invoke and in what order.
@@
Contrary to expectations, recent empirical studies in multi-source RAG show that brute-force, all-in-one retrieval can outperform prompt-based selective routing(Xu et al., 2025). This suggests that naïve prompt-based routing learned from shallow heuristics does not reliably improve the effectiveness and informativeness of retrieved content; instead, it often amplifies systemic inefficiencies and disrupts cross-source coordination, undermining the intended benefits of selectivity. In other words, when routing is weak, selectivity becomes a liability.
@@
Building on this finding, we extend our observation to the behavioral patterns of RL-trained RAG agents. Prior works have applied reinforcement learning (RL) to enhance agents’ ability for multi‑step reasoning with retrieval tools (Zheng et al., 2025). However, we uncover a recurrent failure pattern that we term Tool‑Call Hacking—a domain‑specific form of reward hacking (Skalse et al., 2025). Specifically, the agent learns tool‑calling behaviors that boost the observed reward signal without genuinely improving grounding, which is highly susceptible to Goodhart’s Law(Majka and El-Mhamdi, 2025). We observe two consistent manifestations: (i) Mode collapse via tool overuse: policies default to a narrow subset of tools. (ii) Hallucinated tool use: tool calls are made, but the reasoning chain shows no causal dependence on the retrieved evidence. In both cases, the agent superficially satisfies the reward by producing spurious querying.
@@
To this end, we propose an evidence-grounded agentic reinforcement learning framework, Proof-of-Use (PoU), that explicitly enforces the verifiable linkage between an agent’s reasoning process and the evidence it retrieves. Unlike conventional agentic RL paradigms that optimize for task-level success without scrutinizing how evidence is used, PoU reconstructs the entire agent–environment interaction contract into a stepwise proof schema. Within this schema, each reasoning step must first render an explicit verdict on the usefulness of retrieved evidence and then cite the specific supporting sources—thereby making the evidence–reasoning dependency a learnable and auditable interface.
This contract extends beyond intermediate reasoning: the same evidence–reasoning linkage is propagated to the answer level, where final predictions are judged by their factual consistency with the aggregated cited evidence. Moreover, the training objective itself is grounded in the contract—reward functions are computed on citation validity, perturbation sensitivity, and answer–evidence alignment, while rollout masking is selectively relaxed to preserve truly referenced snippets. Through this tightly coupled design, PoU transforms reasoning supervision from heuristic imitation into a contract-driven optimization of verifiable grounding, enabling the agent to progressively align its internal reasoning dynamics with external factual dependencies.
@@
In this work, we define “tools” as black-box(Ajwani et al., 2024) proxies that interface with external knowledge sources: including but not limited to open-web search, internal document collections, structured APIs, or specialized knowledge bases. This reflects a common abstraction in RAG systems, where tools are used to fetch potentially relevant information conditioned on the current query or reasoning context. While other forms of external tools (e.g., calculators, code interpreters, or simulators) have been explored in agent-based systems(Wölflein et al., 2025), our focus is on the multi-source retrieval setting where the challenge lies in tool routing and evidence integration, not symbolic execution or procedural control. This scoped definition allows us to isolate and study a critical bottleneck in retrieval-augmented reasoning: the agent’s ability to strategically call and utilize retrieval interfaces in a way that causally supports final outputs. We believe this setup is both practically relevant and theoretically rich, and we leave extension to broader tool types as future work.
Contributions.
@@
  • We identify and formalize the tool-call hacking phenomenon of RAG agents in multi-source environments, revealing how sparse rewards and ungrounded rollouts lead to tool overuse and hallucinated grounding.
  • We propose Proof-of-Use (PoU), a contract-driven and evidence-grounded reinforcement learning framework that tightly couples the agent’s interaction protocol with causal evidence–reasoning–answer alignment, yielding verifiable and interpretable reasoning behaviors.
  • Extensive experiments across open-domain and multi-hop QA benchmarks demonstrate that PoU achieves strong performance in accuracy and faithfulness.
@@ 2. Related Works
2.1. Reinforcement Learning based Deepresearch Agent
Large language models (LLMs) are often constrained by the static knowledge embedded in their parameters, making it difficult for them to acquire the latest information in a timely manner. Consequently, when facing time-sensitive or knowledge-intensive questions, they may produce inaccurate answers or hallucinations. Building on this, RAG agents have been extended beyond static retrieval heuristics to more adaptive forms of reasoning and information seeking. Recent research has explored applying reinforcement learning (RL) to enable models autonomously plan reasoning steps, decide when to invoke what retrieval tools, and verify information across different sources (Zheng et al., 2025; Song et al., 2025). For instance, the Search-R1(Jin et al., 2025) framework allows a model to flexibly invoke search engines during answer generation, achieving a closed-loop “reason-and-search” process. The R1-Searcher(Song et al., 2025) method adopts a two-stage RL training approach that significantly enhances autonomous search capability. Similarly, ReSearch(Chen et al., 2025) enables LLMs to treat search as part of the reasoning chain even without supervised reasoning step labels. Furthermore, DeepResearcher(Zheng et al., 2025) extends the training environment to the open Internet, achieving end-to-end reinforcement learning in real-world web settings.
2.2. Reinforcement Learning Algorithms for Agent Training
In terms of optimization algorithms, mainstream approaches primarily adopt policy-gradient methods, such as Proximal Policy Optimization (PPO)(Schulman et al., 2017) and its variants. The Search-R1(Song et al., 2025) conducted a detailed comparison between PPO and the GRPO(Shao et al., 2024) in this setting. The conclusion is that GRPO converges faster, since it reduces variance by using the mean reward of multiple sampled outputs as a baseline. However, in later training stages, GRPO can exhibit reward instability or even collapse. In contrast, PPO yields a smoother and more stable training process, achieving slightly higher final performance. Ultimately, both methods reach comparable reward levels and effectively optimize Search-R1. Differently, R1-Searcher(Song et al., 2025) employs an improved version of the REINFORCE++(Hu et al., 2025) algorithm (a model-free policy gradient approach) combined with KL-divergence regularization and other stabilization techniques. In addition, the previously mentioned loss masking of retrieval results also contributes to stability—by excluding the environment’s textual feedback from gradient computation, the model avoids external noise interfering with its gradients, thus improving convergence and policy consistency. However, in our method, we tightly couple the retrieved evidence, the model’s reasoning process, and the final conclusion through an explicit contract mechanism. Under this formulation, we selectively unmask the returned evidence segments during training, thereby further encouraging the model to achieve grounded alignment between evidence and reasoning and resolve the tool-call hacking.
@@ 2.3. Reward Hacking in Reinforcement Learning
“Reward hacking” (a.k.a. specification gaming)—agents exploiting flaws in objective design to score high reward while failing the designer’s intent—was foregrounded as a core AI-safety problem in Concrete Problems in AI Safety and popularized through empirical case studies and taxonomies of gaming behaviors(Amodei et al., 2016). A formal treatment views reward hacking as arising from misspecified proxies and shows conditions under which “hackability” exists, alongside principled defenses; complementary work analyzes reward tampering, where agents manipulate the reward channel itself(Skalse et al., 2025). These phenomena are closely linked to Goodhart’s Law—over-optimization of a proxy decouples it from the true goal—and to goal misgeneralization, where systems pursue unintended objectives despite correct training feedback, both documented across deep learning and RL settings(Manheim and Garrabrant, 2019).
In post-training for LLMs, reward-model optimization introduces additional failure modes: reward model over-optimization can degrade human-judged quality as policies overfit artifacts of the learned reward. Recent studies propose constrained RLHF and composite-reward formulations to counteract these effects, while iterative RLHF empirically reduces over-optimization over successive rounds but with diminishing returns(Moskovitz et al., 2024). Process-based supervision (e.g., PRMs or step-wise CoT rewards) can provide denser, less gameable signals, though monitoring the reasoning process also creates new attack surfaces where models obfuscate or “game the monitor,” motivating hybrid outcome-plus-process objectives. Orthogonal alignment methods such as Constitutional AI reduce reliance on noisy human labels via principle-guided self-critique; while not a direct fix for reward hacking, they reframe supervision to limit exploitable gaps in reward design(Bai et al., 2022).
Modern RL-driven retrieval–reasoning agents operate in an expanded action space where the policy can call external tools (search engines, browsers, knowledge-base APIs, calculators, etc.) and thereby influence the information source used to compute downstream rewards. Therefore, we notice and formalize a class of prevalence failure modes in this expanded setting as tool-call hacking, and we aim to introduce a solution to this problem in this paper.
@@ 3. Method
A PoU agent follows an iterative reasoning–acting–observing(Yao et al., 2023) trajectory until sufficient information is accumulated to produce a final answer.
3.1. Interaction protocol and proxies
@@ 3.1.2. Proxy Response
Here, the id of each reference is automatically assigned. All four proxies adopt this schema, so the model can cite evidence via <ref>ID_1,ID_2</ref> regardless of the source.
Granularity. For the Web Search Proxy, ref items are organized at the webpage level; typically contain a page title, URL, and snippet. For the Web Browsing, Local Search, and Knowledge Graph proxies, ref items are returned at the paragraph (semantic-chunk) level to enable finer-grained grounding.

### v2

@@ Abstract.
Deep research agents increasingly operate in complex multi-source environments, where large language models interact with external knowledge through iterative tool calls to support multi-step reasoning. While reinforcement learning (RL) enhances their ability to plan and reason across retrieval steps, we identify a critical failure mode in this setting: Tool-Call Hacking. Unlike execution-based tools (e.g., code or math), whose effects are directly observable, the weak observability of causal dependencies between retrieved evidence and reasoning under format- and outcome-level supervision enables agents to maximize surface-level reward signals without genuinely grounding their reasoning in the returned evidence. This leads to distinctive pathologies, including mode collapse via tool overuse and hallucinated tool usage where tool calls are largely decorative.
To address this issue, we propose Proof-of-Use (PoU), an evidence grounded RL framework that explicitly optimizes the causal dependency from retrieval to reasoning and final answers. PoU re-fomulate a fine-grained stepwise interaction protocol in which agents must auditably cite normalized evidence identifiers. We operationalize this via a multi-objective reward design consisting of: (1) two progressive process rewards that constrain citation validity at intermediate steps; (2) a global Answer–Support Alignment reward that enforces consistency between final answers and retrieved evidence; and (3) a curriculum-style adaptive reward mixing mechanism that smoothly transitions agents from dense process supervision to sparse outcome-based objectives. Extensive experiments show the strong performance of PoU and demonstrate the effectiveness in mitigating tool-call hacking. Beyond this, PoU exhibits a notable emergent property: adaptive and robust tool-usage patterns naturally arise under domain and tool shifts, even though PoU does not explicitly optimize for tool adaptation.
@@ 1. Introduction
Figure 1. Tool-distribution entropy over training. At each training step, we compute the entropy of the tool usage distribution as $\tilde{H}=-\sum_{i}p_{i}\log p_{i}\,/\,\log K$, where $p_{i}$ is the proportion of calls to tool $i$ and $K$ is the number of available tools. DeepResearcher exhibits a decreasing entropy, which indicates a tool-call hacking pattern (progressive collapsing toward a narrow tool selection). In contrast, PoU shows higher variance in early training, reflecting exploration, and converges to a stable intermediate entropy value, indicating a learned and task-coordinated tool usage pattern.
Deep research agents(Jin et al., 2025; Song et al., 2025; Zheng et al., 2025) are emerging as a powerful paradigm for complex information-seeking tasks, where large language models interact with external knowledge sources through iterative tool calls to search, retrieve, verify, and synthesize evidence. Unlike single-shot retrieval or static prompting, these agents operate in multi-step environments that require not only reasoning over retrieved information, but also planning when, how, and which tools to invoke to progressively construct reliable answers.
In realistic research settings, no single external knowledge source is sufficient(Ma et al., 2025). Modern agents therefore rely on multiple heterogeneous tools, such as web search engines, knowledge graphs, domain-specific databases, or private document collections, each with distinct coverage, structure, and reliability. These tools are exposed as callable interfaces within agentic frameworks, turning tool selection and sequencing into a central component of the agent’s decision-making process. Consequently, deep research agents must simultaneously solve two tightly coupled challenges: grounding intermediate reasoning in retrieved evidence, and strategically routing tool calls across heterogeneous knowledge sources.
@@
From an operational perspective, deep research agents can be viewed as performing multi-round retrieval-augmented generation (RAG). Contrary to expectations, recent empirical studies show that brute-force, all-in-one retrieval can outperform prompt-based selective routing(Xu et al., 2025). This suggests that naïve prompt-based routing learned from shallow heuristics does not reliably improve retrieval effectiveness; instead, it often amplifies systemic inefficiencies and disrupts cross-source coordination. When routing decisions are not grounded in evidence usage, selectivity becomes a liability rather than an advantage.
@@
Building on these observations, we extend the analysis from routing heuristics to the learning dynamics of RL-trained deep research agents. Prior works have applied reinforcement learning (RL) to enhance agents’ ability to multi‑step reasoning with retrieval tools (Zheng et al., 2025). However, we uncover a recurrent failure pattern that we term Tool‑Call Hacking—a domain‑specific form of reward hacking (Skalse et al., 2025). Specifically, the agent learns tool‑calling behaviors that boost the observed reward signal without genuinely improving grounding, which is highly susceptible to Goodhart’s Law(Majka and El-Mhamdi, 2025). We observe two consistent manifestations (as shown in Figure 1): (i) Mode collapse via tool overuse: policies default to a narrow subset of tools; (ii) Hallucinated tool use: tool calls are made, but the reasoning chain shows no causal dependence on the retrieved evidence. Crucially, both failures arise because existing training signals primarily reward the act of calling tools, rather than the act of using them.
@@
As clarified in Paragraph 1, we focus on retrieval tools, whose impact on outputs is propagated through an implicit, semantic level evidence-reasoning-answer causal chain. In contrast, execution-style tools such as code or mathematical solvers have directly observable contributions, which makes retrieval tools easy to bypass or imitate without genuine informational dependence. To this end, rather than treating evidence grounding as an emergent side effect of task success, we argue that it should be an explicit learning objective in agentic reinforcement learning. We propose Proof-of-Use (PoU), an evidence-grounded framework that refines the coarse-grained supervision commonly used in agentic RL. PoU is built on a stepwise interaction protocol: tool responses are normalized into evidence snippets with stable IDs. Upon observing them, the model must output a usefulness verdict and—if useful—cite the corresponding evidence IDs before reasoning. This exposes a verifiable dependency between retrieved evidence, reasoning, and decisions. On top of this protocol, PoU applies two progressive process rewards to strictly constrain step-wise reasoning. Furthermore, to address the reward sparsity and non-smooth gradients of outcome-only supervision, we introduce a global Answer–Support Alignment reward. This adheres to the intuition that effective learning should optimize the reasoning process alongside the final result. Crucially, we integrate these objectives via a curriculum-style adaptive reward mixing mechanism. This mechanism smoothly transitions focus from dense reasoning signals to final answer correctness, creating a stable gradient ascent path that bridges the gap to sparse outcome-based rewards.
Together, these design choices couple every stage into an auditable Proof-of-Use chain: tool outputs are anchored by stable IDs, steps explicitly cite evidence, cited grounds are selectively perturbed for verification, and the final answer is rewarded for consistency with the same evidence.
@@
In this work, we define “tools” as black-box(Ajwani et al., 2024) proxies that interface with external knowledge sources: retrieval-style tools including but not limited to open-web search, internal document collections, structured APIs, or specialized knowledge bases. This reflects a common abstraction in RAG systems, where tools are used to fetch potentially relevant information conditioned on the current query or reasoning context. We do not extend our analysis to coding or mathematical tools(wölflein2025llmagentsmakingagent).
This choice is deliberate: coding and math tools primarily stress symbolic reasoning and execution correctness, where tool outputs are directly consumed as final answers and failures are typically observable (e.g., wrong values or execution errors). In contrast, retrieval tools serve as intermediate evidence providers. Their effectiveness depends on problem analysis, retrieval planning, and information integration, creating a setting where agents can satisfy tool-call specifications without causally relying on the returned evidence. This makes retrieval-centric tools particularly susceptible to tool-call hacking (see Section LABEL:def-toolcallhack for a formal definition), motivating our PoU rewards.
Contributions:
@@
  • We characterize and analyse a subtle yet under-explored reward-hacking risk tool-call hacking in RL-trained RAG agents under multi-source retrieval settings. We formalize its manifestations and empirically underscores the necessity of tackling this challenge.
  • To mitigate Tool-call Hacking, we propose Proof-of-Use, an evidence-grounded reinforcement learning framework that introduces a stepwise interaction protocol with progressive process rewards, a global answer-support alignment reward and an adaptive reward mixing training strategy.
  • Extensive experiments across multiple datasets demonstrate the strong performance of PoU and its effectiveness in mitigating tool-call hacking. Beyond this, PoU exhibits notable emergent properties: robustness and adaptive tool-usage patterns naturally arise under domain and tool shifts, even though PoU does not explicitly optimize for them.
@@ 2. Related Works
2.1. Reinforcement Learning based Deep Research Agents
Large language models (LLMs) are limited by the static knowledge embedded in their parameters, which often leads to hallucinations when answering time-sensitive or knowledge-intensive queries. To address this limitation, retrieval-augmented generation (RAG) systems have evolved from static heuristics to adaptive reasoning agents. Recent works apply reinforcement learning (RL) to enable LLMs to autonomously plan reasoning steps, decide when and how to invoke retrieval tools, and verify information across heterogeneous sources (Zheng et al., 2025; Song et al., 2025).
Search-R1 (Jin et al., 2025) introduces a closed-loop reason-and-search framework, allowing models to dynamically invoke search engines during answer generation. R1-Searcher (Song et al., 2025) further adopts a two-stage RL training paradigm that substantially improves autonomous search behaviors. ReSearch (Chen et al., 2025) treats search as an integral part of the reasoning chain without relying on supervised intermediate reasoning labels. Beyond controlled environments, DeepResearcher (Zheng et al., 2025) extends RL training to the open Internet, enabling end-to-end optimization in real-world web settings.
@@ 2.2. Reward Hacking in Reinforcement Learning
Reward hacking (a.k.a. specification gaming) refers to agents exploiting misspecified objectives to achieve high reward while violating the designer’s true intent. It was systematically highlighted as a core AI-safety challenge in Concrete Problems in AI Safety and later analyzed through empirical taxonomies of gaming behaviors (Amodei et al., 2016). Formal treatments attribute reward hacking to proxy misspecification, characterizing conditions under which objectives become hackable and proposing principled defenses; related work studies reward tampering, where agents manipulate the reward channel itself (Skalse et al., 2025). These failure modes are closely connected to Goodhart’s Law—over-optimizing a proxy causes it to decouple from the true goal—and to goal misgeneralization, where models pursue unintended objectives despite correct training feedback (Manheim and Garrabrant, 2019).
In LLM post-training, reward-model optimization introduces additional risks: excessive optimization against learned reward models can degrade human-judged quality as policies overfit spurious reward artifacts. Prior work proposes constrained RLHF and composite reward formulations to mitigate such effects, while iterative RLHF can empirically reduce over-optimization across rounds, albeit with diminishing returns (Moskovitz et al., 2024). Process-based supervision, such as process reward models (PRMs) or step-wise Chain-of-Thought(Wei et al., 2023) rewards, offers denser and potentially less gameable signals; however, monitoring intermediate reasoning also opens new attack surfaces, enabling models to obfuscate or game the monitor itself. This has motivated hybrid objectives that jointly supervise both outcomes and processes. Orthogonal alignment approaches, such as Constitutional AI, further reduce reliance on noisy human labels through principle-guided self-critique, reframing supervision to limit exploitable gaps in reward design (Bai et al., 2022).
Modern RL-trained retrieval–reasoning agents further expand the action space by allowing policies to invoke external tools (e.g., search engines, web browsers, or knowledge-base APIs), thereby influencing the information sources used to compute downstream rewards. In this setting, reward hacking manifests in subtle and prevalent forms. We identify and formalize these failures as tool-call hacking, and this paper aims to provide a principled solution to this problem.
@@ 3. Method
PoU follows an iterative reasoning–acting–observing trajectory (Yao et al., 2023) for tool-augmented question answering. To address the often overlooked tool-call hacking in such settings, we design a set of specialized training rewards together with a matched step-level protocol that governs the interaction between reasoning, tool calls, and answers. A key property of PoU is ID-consistent end-to-end coupling: the same evidence IDs introduced by tools are (1) referenced at the step level via usefulness verdicts and citations, and (2) reused at the answer level for evidence–answer consistency checks. This shared identifier layer turns evidence use into a traceable protocol, making every reward attributable to specific retrieved evidence rather than superficial tool calls.
We will first formalize the tool-call hacking phenomenon, and then present the design of PoU rewards in detail.
3.1. Preliminaries
RL Deep Research Agent Interaction Protocol and Rewards.
Traditional deep research agents(Zheng et al., 2025; Song et al., 2025; Jin et al., 2025; Li et al., 2025a) follow a fixed interaction protocol that iterates over <think> $\rightarrow$ <tool_call> $\rightarrow$ <tool_response> in a loop until termination. At each step, the agent observes the input prompt or tool response of the last step, produces an reasoning trace, and then optionally invokes a tool before proceeding to the next step.
Training supervision is typically provided only at the format and answer levels. The overall reward is commonly defined as a weighted aggregation:
(1) $R=\lambda\,R_{\text{format}}+(1-\lambda)\,R_{\text{answer}},$
where $R_{\text{format}}$ enforces syntactic validity of the interaction protocol (e.g., well-formed tags and tool calls), $R_{\text{answer}}$ evaluates the correctness of the final answer, and $\lambda\in[0,1]$ controls the trade-off between format compliance and answer quality.
Tool-Call Hacking (Definition):
We define tool-call hacking as a specific form of reward hacking (Skalse et al., 2025) in tool-augmented LLM agents, where the policy attains high training rewards by satisfying tool-call format requirements and answer-level correctness objectives, while its stepwise reasoning and action behavior does not causally depend on the content returned by the tools. As a result, the trained model deviates from the intended goal of retrieval-augmented generation, introducing a high but often hard-to-detect risk of hallucinated reasoning and answers.
Manifestations of Tool-Call Hacking:
In practice, tool-call hacking exhibits two direct and recurring manifestations in retrieval-augmented agents trained with format- and answer-level rewards.
First, the agent’s tool usage becomes poorly coordinated and collapses to a narrow or repetitive calling pattern. Instead of adaptively planning tool invocations based on the query structure and intermediate findings, the policy tends to overuse a single tool or a fixed call template that reliably satisfies format constraints.
Second, the agent’s stepwise reasoning or rollout-level answers become causally disconnected from the retrieved content. That is, semantically meaningful perturbations to the tool-returned results do not induce corresponding changes in intermediate reasoning steps or the final answer.
Figure 2. The Proof-of-Use Framework
3.2. Settings of PoU
@@ 3.2.2. Proxy Response
Here, the globally unique id of each reference is automatically assigned. All four proxies adopt this schema, so the model can cite evidence via <ref>ID_1,ID_2</ref> regardless of the source.
Granularity. For the Web Search Proxy, ref items are organized at the webpage level, typically including the page title, URL, and a short snippet. In contrast, the Web Browsing, Local Search, and Knowledge Graph proxies return ref items at the paragraph (semantic-chunk) level, enabling finer-grained evidence grounding.